# Universally Unfiltered and Unseen: Input-Agnostic Multimodal Jailbreaks against Text-to-Image Model Safeguards

### v1

@@ Abstract.
Various (text) prompt filters and (image) safety checkers have been implemented to mitigate the misuse of Text-to-Image (T2I) models in creating Not-Safe-For-Work (NSFW) content. In order to expose potential security vulnerabilities of such safeguards, multimodal jailbreaks have been studied. However, existing jailbreaks are limited to prompt-specific and image-specific perturbations, which suffer from poor scalability and time-consuming optimization. To address these limitations, we propose Universally Unfiltered and Unseen (U3)-Attack, a multimodal jailbreak attack method against T2I safeguards. Specifically, U3-Attack optimizes an adversarial patch on the image background to universally bypass safety checkers and optimizes a safe paraphrase set from a sensitive word to universally bypass prompt filters while eliminating redundant computations. Extensive experimental results demonstrate the superiority of our U3-Attack on both open-source and commercial T2I models. For example, on the commercial Runway-inpainting model with both prompt filter and safety checker, our U3-Attack achieves higher success rates than the state-of-the-art multimodal jailbreak attack, MMA-Diffusion. Content Warning: This paper includes examples of NSFW content.
@@ 1. Introduction
To address these limitations, we propose Universally Unfiltered and Unseen (U3)-Attack, a universal multimodal jailbreak attack against T2I models. U3 is Universal, applicable across diverse image inputs and target prompts containing the same sensitive word; Unfiltered, capable of evading prompt filters; and Unseen, capable of evading safety checkers. Specifically, for the image modality attack, we introduce an image-agnostic adversarial patch optimization strategy that synthesizes a universal patch without requiring privileged access to internal model parameters (see image modality in Figure 1 (b)). For the text modality attack, we propose a prompt-agnostic paraphrase set optimization method that derives a reusable paraphrase set from a given sensitive word (see text modality in Figure 1 (b)). This enables the construction of adversarial prompts for multiple target prompts containing the same sensitive word (e.g., ”naked”) using a shared paraphrase set, eliminating the need for repeated optimization. Experimentally, we have effectively explored the security risks of several popular T2I models (SDv1.5, SLD, ESD) as well as commercial online T2I services (Runway-inpainting, Leonardo.Ai, DALL·E 3, and Runway). Our main contributions are as follows:
@@ 3.2. Image-Agnostic Optimization
Safety Checker-Guided Initialization of Adversarial Patch. We adopt the built-in safety checker (SDSC) from Stable Diffusion v1.5 (SDv1.5) (Rombach et al., 2022b) as our target model. Given an input image, the image encoder of SDSC maps it into a latent vector . The SDSC then sequentially calculates the cosine distances between the latent vector and each of the built-in default NSFW concept embeddings, denoted as for = 1, …, . If any cosine distance exceeds the threshold associated with a specific concept embedding, the synthesized image will be flagged as corresponding to that NSFW concept. We aim to craft an adversarial patch that, when applied to an NSFW image —the direct input to SDSC—produces an adversarial image capable of evading SDSC detection. Our objective is formalized as follows:
where denotes the Hadmard product, denotes the cover perturbation that carries the adversarial patch, and denotes a binary mask for used to constrain the location and shape of patch. has the same dimension as which represents a all-one matrix. denotes the latent vector derived from the computation of . is an indicator function that dynamically selects loss terms where the cosine distance exceeds the corresponding threshold. The detailed optimization procedure is outlined in the satge 1 of Figure 3 and Algorithm 1 in Appendix.
Robustness Enhancement of Adversarial Patch. After obtaining the optimal cover perturbation , we apply it to the original input image following Equation 2 to launch the attack. However, although is applied exclusively to the non-inpainted region of , it undergoes subtle variations after passing through the image inpainting model. These variations weaken the adversarial pattern in the corresponding regions of the synthesized output , thereby reducing the overall attack efficacy—an effect we further examine in Section 4.4. To mitigate this issue, we initialize the perturbation variable with (i.e., ), and subsequently perform robustness-oriented fine-tuning. Inspired by physical-world adversarial attacks, where transformations from the data domain to the physical space must be explicitly modeled (Athalye et al., 2018), we propose a residual modeling strategy to characterize the variation in the cover perturbation before and after processing by the inpainting model, which can be formally expressed as:
is a masked image that serves as the image input of model, highlighting the regions that require inpainting. represents the adversarial image, also serving as the input to models. is the text prompt input to the model, which describes the content to be modified and provides guidance for the inpainting process. refers to the synthesized image, which is the model’s output. stands for image inpainting model(e.g., SDv1.5). When calculating the variation , we only need the inputs and outputs of the image inpainting model, without requiring any detail of its internal mechanics.
@@
We ultimately achieve a robust and effective cover perturbation . The specific details are provided in the stage 2 of Figure 3 and Algorithm 2 in Appendix. Notably, throughout all aforementioned processes, the optimization of the cover perturbation , which embeds the adversarial patch, is exclusively directed at the safety checker SDSC. This ensures that gradient updates are propagated solely through its parameters.
@@ 3.3. Prompt-Agnostic Optimization
Construction of paraphrase set. The Latent Diffusion Model (LDM)(Rombach et al., 2022a), instantiated as SDv1.5 (Rombach et al., 2022b), denoises an image within its latent space, with the process guided by text embeddings obtained by encoding the target prompt with the text encoder of CLIP (Radford et al., 2021). Our goal is to ensure that the adversarial prompt does not contain any sensitive word, while still allowing the semantic information associated with the sensitive word to control the final synthesized image. To achieve this, we shift our focus away from the context where the sensitive word appears and instead construct a prompt-agnostic paraphrase set corresponding to each sensitive word. represents the size of the paraphrase set. By ensuring the identical latent features produced by the , given by i.e., , we select the paraphrase for the paraphrase set . We ensure the semantic consistency between the sensitive word and paraphrase by maximizing the cosine similarity between the latent feature and . We formalize the attack objective as follows:
@@
Optimal Paraphrase Selection and Adversarial Prompt Construction. The paraphrase set corresponding to a specific sensitive word remains universally applicable across different target prompts that contain . For a given target prompt containing the sensitive word , we first identify the optimal paraphrase from the corresponding paraphrase set and seamlessly substitute within the target prompt. The criteria for selecting the optimal paraphrase is as follows:
@@ 4.2. Jailbreaking Text-Driven Image Inpainting Models
Bypassing Open Source Safety Checker. To ensure a fair comparison with input-specific attack methods such as MMA and Sneaky, we introduce a variant of our method, denoted as U3-Attack (image-specific), in which a dedicated adversarial patch is generated for each of the 61 test cases. As shown in Table 2, U3-Attack (image-agnostic) achieves an ASR-4-1 of 98.361% on the SDv1.5 inpainting model equipped with the SDSC (CompVis, 2024) under white-box settings, outperforming MMA (83.607%) and Sneaky (59.016%). This performance gain is largely attributed to the flexibility of our patch-based formulation, which allows unconstrained pixel-level perturbations and significantly enhances attack efficacy. Under black-box settings, U3-Attack (image-agnostic) achieves ASR-4-1 scores of 54.098% and 36.066% against the MHSC (Qu et al., 2023) and Q16 (Schramowski et al., 2022), respectively. The strong cross-checker transferability suggests that our adversarial patch captures generalizable adversarial patterns, remaining effective even against safety checkers with architectures entirely dissimilar to SDSC. Compared to image-specific variant, the U3-Attack (image-agnostic) achieves an average improvement of 18 percentage points on SDSC. This is mainly because the image-specific variant is prone to overfitting due to per-image optimization, resulting in reduced robustness. In contrast, the U3-Attack (image-agnostic) benefits from diverse training inputs, leading to more stable patterns and higher success rates. Representative results are shown in the left panel of Figure 6.
@@ 4.3. Jailbreaking Text-to-Image Generation Models
Attack Efficiency. Table 5 highlights the clear efficiency advantage of the text modality attack in U3-Attack, which is primarily due to our streamlined paraphrase selection strategy. For target prompts that share the same sensitive word, we directly select the optimal paraphrase from a predefined paraphrase set, eliminating the need for repeated optimization. In contrast, methods such as QF, MMA, and Sneaky regenerate adversarial prompts from scratch for each target prompt, resulting in significantly higher computational overhead.
@@ 4.4. Ablation Studies
Patch Position. We evaluate four fixed patch positions: top-left (tl), top-right (tr), bottom-left (bl), and bottom-right (br), all chosen for their ample background space to reduce visual saliency and maintain image realism. As shown in Table 7, our adversarial patch consistently achieves high ASR-4-1 across all locations, with success rates exceeding 95%. The top-left position yields the highest rate (98.3%), possibly due to spatial biases in the model or reduced scrutiny by safety mechanisms in that region. These results underscore the robustness and spatial adaptability of our image modality attack in U3-Attack.
@@ 5. Conclusion and Outlook
In this work, we present Universally Unfiltered and Unseen (U3)-Attack, a universal multimodal jailbreak framework against prompt filters and safety checkers in T2I models. Unlike prior methods, U3-Attack is input-agnostic, being applicable to diverse image content and eliminating time-consuming per-prompt optimization. Experiments on state-of-the-art open-source T2I models and commercial APIs demonstrate the superiority of our U3-Attack. We hope that by exposing critical safety flaws in current safeguards, our jailbreak method can be potentially used to help design more robust safeguards. Future work should look like more natural-looking prompts and image perturbations. In addition, extending U3-Attack to text-to-video generators using frame-agnostic perturbations is a promising direction for uncovering temporal vulnerabilities.
@@ B.2. Baselines
Sneaky (Yang et al., 2024c): Sneaky repeatedly interacts with the text-to-image generative model, perturbing the prompt tokens based on query results and utilizing reinforcement learning to guide their perturbation. MMA (Yang et al., 2024a): MMA bypasses prompt filter by generating unconstrained adversarial prompts and evades safety checker by adding imperceptible perturbations to images. QF (Zhuang et al., 2023): QF is an adversarial framework tailored for T2I models. It facilitates the generation of adversarial prompts without requiring direct access to the model, effectively guiding the synthesis process toward images that are semantically misaligned with the original intent.

### v2

@@ Abstract.
Various (text) prompt filters and (image) safety checkers have been implemented to mitigate the misuse of Text-to-Image (T2I) models in creating Not-Safe-For-Work (NSFW) content. In order to expose potential security vulnerabilities of such safeguards, multimodal jailbreaks have been studied. However, existing jailbreaks are limited to prompt-specific and image-specific perturbations, which suffer from poor scalability and time-consuming optimization. To address these limitations, we propose Universally Unfiltered and Unseen (U3)-Attack, a multimodal jailbreak attack method against T2I safeguards. Specifically, U3-Attack optimizes an adversarial patch on the image background to universally bypass safety checkers and optimizes a safe paraphrase set from a sensitive word to universally bypass prompt filters while eliminating redundant computations. Extensive experimental results demonstrate the superiority of our U3-Attack on both open-source and commercial T2I models. For example, on the commercial Runway-inpainting model with both prompt filter and safety checker, our U3-Attack achieves higher success rates than the state-of-the-art multimodal jailbreak attack, MMA-Diffusion. Code is available at https://github.com/yszbb/U3-Attack. Content Warning: This paper includes examples of NSFW content.
@@ 1. Introduction
To address these limitations, we propose Universally Unfiltered and Unseen (U3)-Attack, a universal multimodal jailbreak attack against T2I models. U3 is Universal, applicable across diverse image inputs and text prompts containing the same sensitive word; Unfiltered, capable of evading (text) prompt filters; and Unseen, capable of evading (image) safety checkers. Specifically, for the image modality attack, we introduce an image-agnostic adversarial patch optimization strategy that bypasses safety checkers by synthesizing a universal adversarial patch without relying on access to internal model parameters (see image modality in Figure 1 (b)). For the text modality attack, we propose a prompt-agnostic paraphrase set optimization method that circumvents prompt filters by constructing a reusable paraphrase set for each sensitive word (see text modality in Figure 1 (b)). This design allows adversarial prompts to be efficiently generated across multiple target prompts containing the same sensitive term (e.g., “naked”), eliminating the need for repeated optimization. Experimentally, we have effectively explored the security risks of several popular T2I models (SDv1.5, SLD, ESD) as well as commercial online T2I services (Runway-inpainting, Leonardo.Ai, DALL·E 3, and Runway). Our main contributions are as follows:
@@ 3.2. Image-Agnostic Optimization
Stage 1: Safety Checker-Guided Initialization of Adversarial Patch. We adopt the built-in safety checker (SDSC) from Stable Diffusion v1.5 (SDv1.5) (Rombach et al., 2022b) as our target model. Given an input image, the image encoder of SDSC maps it into a latent vector . The SDSC then sequentially calculates the cosine distances between the latent vector and each of the built-in default NSFW concept embeddings, denoted as for = 1, …, . If any cosine distance exceeds the threshold associated with a specific concept embedding, the synthesized image will be flagged as corresponding to that NSFW concept. We aim to craft an adversarial patch that, when applied to an NSFW image —the direct input to SDSC—produces an adversarial image capable of evading SDSC detection. Our objective is formalized as follows:
where denotes the Hadmard product, denotes the cover perturbation that carries the adversarial patch, and denotes a binary mask for used to constrain the location and shape of patch. has the same dimension as which represents a all-one matrix. denotes the latent vector derived from the computation of . is an indicator function that dynamically selects loss terms where the cosine distance exceeds the corresponding threshold. The detailed optimization procedure is outlined in Stage 1 of Figure 3 and Algorithm 1 in the Appendix.
Stage 2: Robustness Enhancement of Adversarial Patch. After obtaining the optimal cover perturbation , we apply it to the original input image following Equation 2 to launch the attack. However, although is applied exclusively to the non-inpainted region of , it undergoes subtle variations after passing through the image inpainting model. These variations weaken the adversarial pattern in the corresponding regions of the synthesized output , thereby reducing the overall attack efficacy—an effect we further examine in Section 4.4. To mitigate this issue, we initialize the perturbation variable with (i.e., ), and subsequently perform robustness-oriented fine-tuning. Inspired by physical-world adversarial attacks, where transformations from the data domain to the physical space must be explicitly modeled (Athalye et al., 2018), we propose a residual modeling strategy to characterize the variation in the cover perturbation before and after processing by the inpainting model, which can be formally expressed as:
is a masked image that serves as the image input of model, highlighting the regions that require inpainting. represents the adversarial image, also serving as the input to models. is the text prompt input to the model, which describes the content to be modified and provides guidance for the inpainting process. refers to the synthesized image, which is the model’s output. stands for image inpainting model (e.g., SDv1.5). When calculating the variation , we only need the inputs and outputs of the image inpainting model, without requiring any details of its internal mechanics.
@@
We ultimately achieve a robust and effective cover perturbation . The specific details are provided in Stage 2 of Figure 3 and Algorithm 2 in the Appendix. Notably, throughout all aforementioned processes, the optimization of the cover perturbation , which embeds the adversarial patch, is exclusively directed at the safety checker SDSC. This ensures that gradient updates are propagated solely through its parameters.
@@ 3.3. Prompt-Agnostic Optimization
Construction of paraphrase set. The Latent Diffusion Model (LDM)(Rombach et al., 2022a), instantiated as SDv1.5 (Rombach et al., 2022b), denoises an image within its latent space, with the process guided by text embeddings obtained by encoding the target prompt with the text encoder of CLIP (Radford et al., 2021). Our goal is to ensure that the adversarial prompt does not contain any sensitive words, while still allowing the semantic information associated with the sensitive word to control the final synthesized image. To achieve this, we shift our focus away from the context where the sensitive word appears and instead construct a prompt-agnostic paraphrase set corresponding to each sensitive word. represents the size of the paraphrase set. By ensuring the identical latent features produced by the , given by i.e., , we select the paraphrase for the paraphrase set . We ensure the semantic consistency between the sensitive word and paraphrase by maximizing the cosine similarity between the latent feature and . We formalize the attack objective as follows:
@@
Optimal Paraphrase Selection and Adversarial Prompt Construction. The paraphrase set corresponding to a specific sensitive word remains universally applicable across different target prompts that contain . For a given target prompt containing the sensitive word , we first identify the optimal paraphrase from the corresponding paraphrase set and seamlessly substitute within the target prompt. The criterion for selecting the optimal paraphrase is as follows:
@@ 4.2. Jailbreaking Text-Driven Image Inpainting Models
Bypassing Open Source Safety Checker. To ensure a fair comparison with input-specific attack methods such as MMA and Sneaky, we introduce a variant of our method, denoted as U3-Attack (image-specific), in which a dedicated adversarial patch is generated for each of the 61 test cases. As shown in Table 2, U3-Attack (image-agnostic) achieves an ASR-4-1 of 98.361% on the SDv1.5 inpainting model equipped with the SDSC (CompVis, 2024) under white-box settings, outperforming MMA (83.607%) and Sneaky (59.016%). This performance gain is largely attributed to the flexibility of our patch-based formulation, which allows unconstrained pixel-level perturbations and significantly enhances attack efficacy. Under black-box settings, U3-Attack (image-agnostic) achieves ASR-4-1 scores of 54.098% and 36.066% against the MHSC (Qu et al., 2023) and Q16 (Schramowski et al., 2022), respectively. The strong cross-checker transferability suggests that our adversarial patch captures generalizable adversarial patterns, remaining effective even against safety checkers with architectures entirely dissimilar to SDSC. Compared to the image-specific variant, the U3-Attack (image-agnostic) achieves an average improvement of 18 percentage points on SDSC. This is mainly because the image-specific variant is prone to overfitting due to per-image optimization, resulting in reduced robustness. In contrast, the U3-Attack (image-agnostic) benefits from diverse training inputs, leading to more stable patterns and higher success rates. Representative results are shown in the left panel of Figure 6.
@@ 4.3. Jailbreaking Text-to-Image Generation Models
Attack Efficiency. Table 5 highlights the efficiency advantage of the text modality attack in U3-Attack, which is primarily due to our streamlined paraphrase selection strategy. For target prompts that share the same sensitive word, we directly select the optimal paraphrase from a predefined paraphrase set, eliminating the need for repeated optimization. In contrast, methods such as QF, MMA, and Sneaky regenerate adversarial prompts from scratch for each target prompt, resulting in significantly higher computational overhead.
@@ 4.4. Ablation Studies
Patch Position. We evaluate four fixed patch positions: top-left (tl), top-right (tr), bottom-left (bl), and bottom-right (br), all chosen for their ample background space to reduce visual saliency and maintain image realism. As shown in Table 7, our adversarial patch consistently achieves high ASR-4-1 across all locations, with success rates exceeding 95%. The top-left position yields the highest rate (98.3%), possibly due to spatial biases in the model or reduced scrutiny by safety mechanisms in that region. These results underscore the robustness of our U3-Attack.
@@ 5. Conclusion and Outlook
In this work, we present Universally Unfiltered and Unseen (U3)-Attack, a universal multimodal jailbreak framework against prompt filters and safety checkers in T2I models. Unlike prior methods, U3-Attack is input-agnostic, being applicable to diverse image content and eliminating time-consuming per-prompt optimization. Experiments on state-of-the-art open-source T2I models and commercial APIs demonstrate the superiority of our U3-Attack. We hope that by exposing critical safety flaws in current safeguards, our jailbreak method can be potentially used to help design more robust safeguards. Future work should look at more natural-looking prompts and image perturbations. In addition, extending U3-Attack to text-to-video generators using frame-agnostic perturbations is a promising direction for uncovering temporal vulnerabilities.
@@ B.2. Baselines
Sneaky (Yang et al., 2024c): Sneaky repeatedly interacts with the text-to-image generative model, perturbing the prompt tokens based on query results and utilizing reinforcement learning to guide their perturbation. MMA (Yang et al., 2024a): MMA bypasses prompt filters by generating unconstrained adversarial prompts and evades safety checkers by adding imperceptible perturbations to images. QF (Zhuang et al., 2023): QF is an adversarial framework tailored for T2I models. It facilitates the generation of adversarial prompts without requiring direct access to the model, effectively guiding the synthesis process toward images that are semantically misaligned with the original intent.